If the product of any two rational numbers is 2 and one of them is 1/7, find the other?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider 2 rational numbers as “a” and “b”.

Given, a = 1/7 and a × b = 2

Now, 1/7 × b = 2

⇒ b = 7 × 2 = 14

So, the other rational number will be 14.

Thanks